Stability of a crystal is reflected in the magnitude of its melting points'. Comment. Collect melting points of solid water, ethyl alcohol, diethyl ether and methane from a data book. What can you say about the intermolecular forces between these molecules?

Grade:8

1 Answers

Gaurav
9 years ago
Helllo Student
Melting points of substances
Water=273K
Ethyl alcohol=155·7K
Diethyl ether=156·8K
Methane=90·5K.
The stability of crystals depends up on the packing of the atoms. if there are strong force of attraction between the constituent particles then greater the stability.

Water– strong intermolecular hydrogen bonding
Ethyl alcohol– strong intermolecular hydrogen bonding.
Diethyl ether– it is a polar molecule so there is dipole –dipole attraction
Methane– isa non–polar molecule so that there is weak van der Waal's forces (London dispersion forces).